# RealRep: Generalized SDR-to-HDR Conversion with Style Disentangled Representation Learning

### v1

@@ 3.2.2. Detail Refinement
Table 1. Comparison of quantitative results on HDRTV4K (known degradation benchmark). All methods were trained using all degraded SDR versions from the HDRTV4K dataset. Bold and underline indicate the best and second-best results, respectively.
@@
Table 2. Comparison of quantitative results on HDRTV1K (unknown degradation benchmark). All methods were trained using all degraded SDR versions from the HDRTV4K dataset. Bold and underline indicate the best and second-best results, respectively.
3.3. Training Strategy
Our training strategy is tailored for the multi-view degradation encoder and the Degradation-Domain Aware Controlled Mapping network. Inspired by MoCo (He et al., 2019), the encoder is paired with a negative network whose weights are updated using Exponential Moving Average (EMA). This negative encoder helps maintain a slowly evolving representation space, providing stable and consistent targets for contrastive learning. The training process is divided into two stages, with specific adjustments in each stage to optimize learning.
In the first stage, contrastive learning is disabled, and the negative networks in the encoder are frozen to avoid premature noise injection. The control convolutions in the modulation modules are initialized as standard convolutions, and modulation is fully enabled. This allows the encoder and backbone to learn stable feature representations and basic degradation-aware modulation, serving as a strong initialization for the second stage. In the second stage, contrastive learning is introduced to improve the encoder’s discrimination of degradation patterns. The control convolutions are re-initialized as zero-initialized layers (ZeroConv), and gradually trained from zero to softly inject degradation priors. This progressive activation stabilizes the training and prevents disruption from strong contrastive priors, enabling effective adaptation to diverse degradations.
The pixel loss function is defined as:
(7) $L_{\text{pix}}=\|I_{\text{HDR}}^{\text{pred}}-I_{\text{HDR}}^{\text{G.T.}}\|_{1}+\lambda_{l}\|I_{\text{GM}}-I_{\text{HDR}}^{\text{G.T.}}\|_{1},$
where $I_{\text{HDR}}^{\text{pred}}$ is the predicted HDR image, $I_{\text{HDR}}^{\text{G.T.}}$ is the ground truth HDR image, and $\lambda_{l}$ is a hyper-parameter that balances the contributions of the global mapping (GM) loss and the final HDR prediction loss. Here, $I_{\text{GM}}$ denotes the intermediate output from the multi-layer DCM module in DDACMNet, which captures the globally modulated representation before detail refinement. Empirically, $\lambda_{l}$ is set to 0.7.
The total loss function integrates the pixel loss and contrastive loss as:
(8) $L_{\text{total}}=L_{\text{pix}}+\lambda_{\text{contra}}L_{\text{contra}}$
where $L_{\text{contra}}$ represents the contrastive loss for robust multi-view representation learning, and $\lambda_{\text{contra}}$ is empirically set to 0.2 to balance the contributions of the two loss terms.
Figure 6. Comparison of t-SNE visualizations for PromptIR (left) and our method (right). Each 2D coordinate represents the projection of a single frame’s feature vector from the corresponding dataset, with PCA (16 components) followed by t-SNE (perplexity=50, n_components=2). The visualizations highlight the structural differences between datasets, with the red points representing features from unknown degradations. This demonstrates our method’s ability to produce a more compact and distinct feature space, as shown by the separation and clustering of both known and unknown degradation types.
4. Experiments
4.1. Experimental Setup
4.1.1. Dataset.
We evaluate the effectiveness of our proposed framework on the HDRTV4K dataset (Guo et al., 2023), which consists of 3878 4K HDR images encoded in BT.2020/PQ1000. All degradation types included in HDRTV4K are utilized for training and testing. The training set contains 3478 images from HDRTV4K, while the testing set comprises 400 images from HDRTV4K. Additionally, we include 117 images from the HDRTV1K dataset (Chen et al., 2021d) in the testing set to evaluate the performance on unknown degradations, resulting in a total of 517 test images.
4.1.2. Implementation Details.
@@ 4.1. Experimental Setup
4.2. Experimental Results
Figure 7. Visual comparison with the state-of-the-art methods. Following hdrtvdm (Guo et al., 2023), the HDR images are shown in their original encoding without tone mapping to preserve highlights. Transfer function and gamut mismatches may cause darker, less saturated grayish appearance on paper or SDR screens.
4.2.1. Quantitative Results.
We validate RealRep on the HDRTV4K (known degradation benchmark) and HDRTV1K (unknown degradation benchmark) datasets, comparing it to state-of-the-art SDR-to-HDR methods, including HDRUNet(Chen et al., 2021a), HDRTVNet(Chen et al., 2021d), FMNet(Xu et al., 2022), ICTCPNet(Huang et al., 2023), LSNet(Guo et al., 2023), PromptIR(Potlapalli et al., 2023), and RAM-PromptIR(Qin et al., 2024).
@@ 4.2. Experimental Results
4.2.2. Qualitative Results.
Figure 7 compares HDR/WCG images generated by state-of-the-art methods and our proposed RealRep framework. Existing methods, such as HDRTVNet, FMNet, and ICTCPNet, struggle with diverse degradations, causing artifacts like banding, color distortion, and uneven luminance due to their reliance on fixed tone mapping strategies, which limits adaptability and leads to inconsistent performance. For instance, HDRTVNet produces overexposure and banding, FMNet shows luminance and chrominance mismatches, and ICTCPNet renders desaturated colors. In contrast, RealRep combines disentangled multi-view degradation representation learning with a domain-adaptive hierarchical transfer framework to dynamically adapt to varying degradations, producing HDR images that are natural, consistent, and artifact-free. These qualities are evident in Figure 6, where RealRep achieves smoother luminance transitions, better highlight recovery, and accurate color distributions, surpassing other methods in addressing unknown degradations and delivering results closer to the ground truth.
4.2.3. Joint Luma-/Chroma-aware Representation Visualization.
@@
Table 3. Ablation on multi-view representations and multi-view fusion, evaluated by PSNR (dB). The final column reports the average across all eight degradation types. Due to space limitations, only four representative degradations are shown.
@@
Table 4. Ablation on the degradation-aware controlled mapping module, evaluated by PSNR (dB).
4.3. Ablation Studies.
@@
In this section, we conduct a series of ablation experiments to investigate the effectiveness of the decoupled multi-view representations and degradation-aware controlled mapping module.
4.3.1. Ablation of Disentangled Multi-View Representations
@@
To evaluate the effectiveness of disentangled multi-view representations, we conduct ablation experiments on the luminance representation $z_{\text{lum}}$, chrominance representation $z_{\text{chr}}$, their combination, the contrastive loss $\mathcal{L}_{\text{contra}}$, and the multi-view fusion module. As shown in Table 3, the baseline model, which lacks all multi-view representations and fusion, achieves the lowest average PSNR of 25.74dB, highlighting the necessity of these components. Removing either $z_{\text{lum}}$ or $z_{\text{chr}}$ reduces the average PSNR to 28.96dB and 29.58dB, respectively, underscoring their independent contributions. The absence of the multi-view fusion module, which integrates luminance and chrominance features, further lowers the average PSNR to 29.83dB, demonstrating its importance for balancing and integrating these representations. Furthermore, removing the contrastive loss $\mathcal{L}_{\text{contra}}$ leads to a performance drop from 31.42dB to 30.04dB, suggesting that the contrastive learning space plays a crucial role in guiding the model to extract informative and discriminative priors. The full RealRep model achieves the highest average PSNR of 31.42dB, validating the joint effectiveness of multi-view representations, fusion, and contrastive supervision in enhancing SDR-to-HDR performance.
4.3.2. Ablation of Degradation-Domain Aware Controlled Mapping Network
@@
We further investigate the impact of the degradation-aware controlled module by conducting ablation experiments on its global and local mapping components, as well as the control mechanism. As shown in Table 4, removing either the global mapping or local mapping leads to a noticeable performance drop, with average PSNRs of 25.00dB and 30.28dB, respectively, validating their complementary roles in capturing hierarchical domain knowledge. Notably, the control mechanism—comprising a zero-initialized convolution and a two-stage training strategy—plays a key role in enhancing generalization. Removing it leads to a decline in both overall performance and generalization ability, with the average PSNR dropping from 31.16dB to 30.78dB, and the PSNR on the challenging HDRTV1K dataset decreasing from 29.12dB to 28.87dB. These results highlight the effectiveness of the control design in improving transferability and robustness across diverse and unseen degradation scenarios. The full RealRep model consistently achieves the best results, demonstrating the importance of the complete degradation-aware controlled mapping design.
4.3.3. Ablation of encoder architecture
We compare ResNet and our UNet-based encoder under identical training settings (batch size = 16, 100K iterations). As shown in Table 5, the UNet encoder achieves better performance (27.28 dB vs. 26.87 dB) with fewer parameters (0.50M vs. 1.23M), demonstrating its superior efficiency and representation capability.
Encoder Parameters (M) PSNR (dB)
ResNet Encoder 1.2289 26.87
UNet Encoder (Ours) 0.5044 27.28
Table 5. Ablation on encoder architecture. UNet-based encoder outperforms ResNet under identical training settings.
5. Concluding Remarks
@@
This paper presents RealRep, a novel SDR-to-HDR method designed for real-world scenarios with diverse SDR styles. RealRep tackles luminance and chrominance variations using style disentangled representation learning, capturing intrinsic style differences to provide realistic guidance for inverse tone mapping. A Degradation-Domain Aware Controlled Mapping framework further enables dynamic luminance range expansion and color gamut extension. Evaluations on diverse and unknown degradations demonstrate our superiority over state-of-the-art methods in visual quality and generalization. Future work will explore extending RealRep to video HDR conversion and real-time applications.

### v2

@@ Introduction
Table 1: Comparison of quantitative results on HDRTV4K (known degradation benchmark). The first group contains CNN-based methods, while the second group includes transformer-based methods. All methods were trained using all degraded SDR versions from the HDRTV4K dataset. Bold and underline indicate the best and second-best results, respectively.
Degradation-Domain Aware Controlled Mapping Network
Degradation-Aware Feature Modulation
To handle spatially-varying SDR degradations, we design a hierarchical degradation-aware feature modulation mechanism composed of two sequential stages: a stack of NN Dense Controlled Mapping (DCM) modules followed by NN Sparse Controlled Mapping (SCM) modules. These modules inject degradation priors into the network via affine transformations over layer-normalized features, enabling both global and local adaptation to degradation patterns. Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), we insert zero-initialized control layers before each modulation to gradually introduce prior information in a stable manner.
Each DCM module performs global, channel-wise modulation conditioned on a compact degradation embedding $z_{\text{deg}}^{g}\in\mathbb{R}^{C\times 1\times 1}$. A pair of zero-initialized $1\times 1$ convolutions predict affine parameters $\gamma$ and $\beta$, which are applied to normalized features as $\hat{x}=\gamma\cdot x+\beta$. To enhance modulation flexibility, we extend this to degradation-aware LayerNorm, where the affine parameters are dynamically predicted from $z_{\text{deg}}^{g}$. We stack NN such DCM modules to progressively inject global degradation priors.
After global modulation, we apply NN SCM modules to further refine features under spatially-varying degradations. Each SCM module utilizes a local degradation map $z_{\text{deg}}^{l}\in\mathbb{R}^{C\times H\times W}$ to generate pixel-wise affine parameters. These are obtained via two parallel branches, each consisting of a shared $1\times 1$ convolution, a ReLU activation, and a zero-initialized $3\times 3$ convolution. The resulting parameters modulate the features using $\hat{x}=\gamma\odot x+\beta$, enabling spatially adaptive transformation. Outputs from both DCM and SCM stages are fused via residual connections, forming a unified degradation-aware modulation pipeline.
Method HDR-VDP3↑ EHL↓ FHLP↓ EWG↓ FWGP↓
HDRTVNet 8.69 0.0079 0.0630 0.0024 0.0287
FMNet 8.75 0.0074 0.0621 0.0022 0.0273
ICTCPNet 8.90 0.0067 0.0580 0.0019 0.0250
LSNet 9.19 0.0060 0.0567 0.0017 0.0241
AirNet 8.48 0.0086 0.0662 0.0027 0.0303
PromptIR 9.11 0.0078 0.0735 0.0012 0.0345
RAM-PromptIR 9.17 0.0070 0.0716 0.0013 0.0329
Ours 9.35 0.0032 0.0420 0.0013 0.0087
Table 2: Quantitative comparison using HDR-specific perceptual quality metrics. $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better. Bold and underline indicate the best and second-best results, respectively.
Detail Refinement
To enhance textures and remove residual artifacts, we append a detail refinement module consisting of stacked ResBlocks:
$I_{\text{HDR}}=F_{\text{refine}}(I_{\text{SCM}})=\text{Conv}_{3\times 3}\circ[\text{ResBlock}]^{N_{r}}\circ\text{Conv}_{3\times 3}(I_{\text{SCM}})$
where $I_{\text{SCM}}$ is the output after the final SCM module, and $N_{r}$ is the number of residual blocks.
Training Strategy
We adopt a two-stage training scheme. In Stage 1, contrastive learning is disabled and control layers are initialized as standard convolutions, allowing stable optimization of the encoder and modulation modules. In Stage 2, contrastive supervision is introduced via an EMA-updated negative encoder (He et al. 2019), and all control layers are re-initialized as zero-convs to softly inject degradation priors.
The total loss combines pixel-level supervision and contrastive regularization:
$\mathcal{L}_{\text{total}}=\|I_{\text{HDR}}^{\text{pred}}-I_{\text{HDR}}^{\text{G.T.}}\|_{1}+\lambda_{l}\|I_{\text{DCM}}-I_{\text{HDR}}^{\text{G.T.}}\|_{1}+\lambda_{\text{contra}}\mathcal{L}_{\text{contra}}$
where $I_{\text{DCM}}$ denotes the output of the final DCM module. We empirically set $\lambda_{l}=0.7$, $\lambda_{\text{contra}}=0.2$.
@@
Table 3: Comparison of quantitative results on HDRTV1K (unknown degradation benchmark). All methods were trained using all degraded SDR versions from the HDRTV4K dataset. Bold and underline indicate the best and second-best results, respectively.
Figure 5: Visual comparison with the state-of-the-art methods. Following hdrtvdm (Guo et al. 2023), the HDR images are shown in their original encoding without tone mapping to preserve highlights. Our method delivers visually superior results, with clearer highlight recovery, more vivid colors, and better structure preservation compared to prior methods. Visualization is best viewed on an HDR screen.
Experiments
Experimental Setup
Dataset.
We evaluate the effectiveness of our proposed framework on the HDRTV4K dataset (Guo et al. 2023), which consists of 3878 4K HDR images encoded in BT.2020/PQ1000. All degradation types included in HDRTV4K are utilized for training and testing. The training set contains 3478 images from HDRTV4K, while the testing set comprises 400 images from HDRTV4K. Additionally, we include 117 images from the HDRTV1K dataset (BT.2020/PQ1000) (Chen et al. 2021d) in the testing set to evaluate the performance on unknown degradations, resulting in a total of 517 test images. The SDR content is configured with BT.709 gamut, gamma 2.2, and a peak brightness of 100 nits.
Implementation Details.
@@
Experimental Results
Quantitative Results.
We validate RealRep on the HDRTV4K (known degradation benchmark) and HDRTV1K (unknown degradation benchmark) datasets, comparing it to state-of-the-art SDR-to-HDR methods, including HDRUNet (Chen et al. 2021c), HDRTVNet (Chen et al. 2021d), FMNet (Xu et al. 2022), ICTCPNet (Huang et al. 2023), LSNet (Guo et al. 2023), PromptIR (Potlapalli et al. 2023), and RAM-PromptIR (Qin et al. 2024). To comprehensively evaluate visual quality, we report both traditional metrics such as PSNR and SSIM, as well as HDR-specific perceptual quality metrics, including HDR-VDP3 and the error-aware HDR metrics EHL, FHLP, EWG, and FWGP proposed by  (Guo et al. 2023). Additionally, we evaluate perceptual color difference using $\Delta E_{\mathrm{ITP}}$, which is reported in the supplementary material.
@@
Qualitative Results.
Figure 5 compares HDR/WCG images generated by state-of-the-art methods and our proposed RealRep framework. Existing methods, such as HDRTVNet, FMNet, and ICTCPNet, struggle with diverse degradations, causing artifacts like banding, color distortion, and uneven luminance due to their reliance on fixed tone mapping strategies, which limits adaptability and leads to inconsistent performance. For instance, HDRTVNet produces overexposure and banding, FMNet shows luminance and chrominance mismatches, and ICTCPNet renders desaturated colors. In contrast, RealRep combines disentangled multi-view degradation representation learning with a domain-adaptive hierarchical transfer framework to dynamically adapt to varying degradations, producing HDR images that are natural, consistent, and artifact-free. These qualities are evident in Figure 5, where RealRep achieves smoother luminance transitions, better highlight recovery, and accurate color distributions, surpassing other methods in handling unknown degradations and producing results closer to the ground truth.
Figure 6: Comparison of t-SNE visualizations for PromptIR (left) and our method (right). Each point represents a frame-level feature projected via PCA (16 components) and t-SNE (perplexity=50). Red points indicate unknown degradations. Our method yields a more compact and well-separated feature space, clearly distinguishing known and unknown degradations.
Joint Luma-/Chroma-aware Representation Visualization.
@@
Table 4: Ablation on multi-view representations and multi-view fusion, evaluated by PSNR (dB). The final column reports the average across all eight degradation types. Due to space limitations, only four representative degradations are shown.
@@
Table 5: Ablation on the degradation-aware controlled mapping module, evaluated by PSNR (dB).
Ablation Studies.
@@
In this section, we perform ablation studies to evaluate the impact of multi-view representations and the degradation-aware controlled mapping module.
Ablation of Disentangled Multi-View Representations
@@
We evaluate the impact of disentangled multi-view representations through ablations on the luminance representation $z_{\text{lum}}$, chrominance representation $z_{\text{chr}}$, their fusion, the contrastive loss $\mathcal{L}_{\text{contra}}$, and the fusion module. As shown in Table 4, the baseline without multi-view components yields the lowest PSNR (25.74dB), confirming their necessity. Removing $z_{\text{lum}}$ or $z_{\text{chr}}$ reduces PSNR to 28.96dB and 29.58dB, respectively, indicating their individual contributions. Without the fusion module, PSNR drops to 29.83dB, showing its role in integrating luminance and chrominance features. Excluding the contrastive loss $\mathcal{L}_{\text{contra}}$ leads to a notable drop from 31.42dB to 30.04dB, highlighting the benefit of contrastive supervision. The full RealRep model achieves the highest PSNR (31.42dB), validating the joint effectiveness of all components in enhancing SDR-to-HDR performance.
Ablation of Degradation-Domain Aware Controlled Mapping Network
@@
We conduct ablation studies on the proposed degradation-aware controlled module, focusing on its global/local mapping components and the control mechanism. As shown in Table 5, removing either the global or local mapping leads to notable performance drops (25.00dB and 30.28dB average PSNR, respectively), confirming their complementary roles in modeling hierarchical domain knowledge. The control mechanism, composed of a zero-initialized convolution and a two-stage training strategy, also proves essential. Its removal reduces average PSNR from 31.16dB to 30.78dB, and performance on HDRTV1K from 29.12dB to 28.87dB, indicating weakened generalization. Overall, the complete RealRep model achieves the best results, demonstrating the effectiveness of the full degradation-aware controlled mapping design.
Effectiveness on Unknown TM-Style SDR Inputs
Our method is designed to generalize better to unknown tone-mapped SDR inputs. As shown in Table 6, our method significantly outperforms a classification-based baseline (ResNet+LSNet) on HDRTV1K, achieving higher PSNR and HDR-VDP3 scores.
Method PSNR HDR-VDP3
ResNet+LSNet 21.57 9.24
Ours 29.12 9.60
Table 6: Generalization to unknown TM-style SDR inputs on HDRTV1K.
Concluding Remarks
@@
This paper introduces RealRep, a novel SDR-to-HDR framework specifically designed to handle the diverse styles and degradations encountered in real-world SDR content. RealRep tackles luminance and chrominance variations using style disentangled representation learning, capturing intrinsic style differences to provide realistic guidance for inverse tone mapping. A Degradation-Domain Aware Controlled Mapping framework further enables dynamic luminance range expansion and color gamut extension. Extensive evaluations on diverse and previously unknown degradations demonstrate that RealRep consistently outperforms state-of-the-art methods in both visual fidelity and generalization capability. Future work will explore extending RealRep to video HDR conversion and real-time applications.